1. If the Center of a circle is at (2, -7) and a point on the circle (5,6) find the formula of the circle.

2. What surfaces in R3 are represented by the following equations?
z = 3
y = 5


3. Find an equation of a sphere with radius r and center C(h, k, l).

4. Show that x2 + y2 + z2 + 4x – 6y + 2z + 6 = 0 is the equation of a sphere. Also, find its center and radius.
1
Expert's answer
2020-10-12T17:45:08-0400

1.The equation of the circle


"(x-x_C)^2+(y-y_C)^2=r^2"

"x_C=2, y_C=-7"

Point "(5,6)" lies on the circle


"(5-2)^2+(6+7)^2=r^2"

"r^2=178"

The equation of the circle


"(x-2)^2+(y+7)^2=178"


2. In three-space, the equation "z=3" represents a horizontal plane, parallel to the xy-plane and at height 3. 

In three-space, the equation "y=5" represents a vertical plane parallel to the xz-plane and passing through the point "(0,5,0)".

3. The general equation of a sphere is:


"(x-h)^2+(y-k)^2+(z-l)^2=r^2 ,"

where "(h,k,l)" represents the center "C" of the sphere, "r" represents the radius, and "x,y," and "z" are the coordinates of the points on the surface of the sphere.

4.


"x^2+y^2+z^2+4x-6y+2z+6=0"

"x^2+4x+4-4+y^2-6y+9-9+z^2+2z+1-1+6=0"

"(x+2)^2+(y-3)^2+(z+1)^2=8"

This is the equation of a sphere with radius "r=2\sqrt{2}" and center "C(-2, 3, -1)."